# A possible statistics loophole in Bell’s theorem

### v2

@@ V.1 Actual upper bound of $S_{N}$
Let us analyze the problem from a practical point of view, considering only the measured data at our disposal [28, 29, 30, 31, 2]. To measure $S_{N}$ we must perform $4N$ experiments, that is we must generate $4N$ distinct photon pairs. After we had ran the experiment $4N$ times using four different pairs of polarization analyzers settings for Alice and Bob, at the end of the day we got an array of $4N$ values, say $V_{N}$, given by
@@
By definition, the elements of the $N\times 4$ matrix $V_{N}^{\text{HV}}$ are equal to either $+1/N$ or $-1/N$, and given by the corresponding elements of $V_{N}$ in (88). If we sum the $\mu^{\text{th}}$ column of $V_{N}^{\text{HV}}$, we obtain the correlation function (84). Clearly, rearranging the order of the elements in each column will not alter the value of such correlation function. So, we can exploit this freedom to recast $V_{N}^{\text{HV}}$ in the new matrix $W_{N}^{\text{HV}}$, defined by
@@ VI Conclusions
In summary, we have demonstrated two things in this paper. First, no counterfactual reasoning is required to prove the CHSH inequality. Second, the upper bound of this inequality can be greater than $2$, but the numerical value of the actual bound is not computable because of our unavoidable ignorance about the probability distribution of the hidden variables. To be more specific, the Bell parameter $S$ is bounded by $\left|S\right|\leq 2$ for HV theories, where $S$ is the mean value of a linear combination of correlation functions:
@@
One final note: given the experimental data set collected, for example, in the form of the matrix $V_{N}$ in (88), a researcher can, of course, perform a full statistical analysis of these data and also infer the rate of convergence of the limit $S_{N}\to S$. However, the moment he or she wants to compare these data with a given physical model, he or she is faced with an obvious imbalance. In fact, while QM provides a complete model that makes it possible to calculate not only the mean values of certain measurable quantities but also their probability distributions, generic hidden-variable theories provide, in practice, only mean values, which are insufficient to infer all the statistical properties of the hidden-variable model. The symmetry between quantum and hidden-variable theories of our statistical analysis presented in Sec. IV, breaks down in the moment the explicit knowledge of the correlation functions $C(\alpha,\beta)$ becomes necessary.
So, in the end, what the researcher can practically do is to verify the agreement of the experimental data with the predictions of quantum mechanics, as after all is done in all meaningful quantum physics experiments (except for those checking Bell’s theorem!) [32]. Because of the relevance of Bell’s theorem in practical applications of quantum mechanics, such as quantum computing or quantum cryptography, we believe that these results have the potential to open up new avenues on this important topic.
@@ A.1 Single experiment
where $k_{\mu},l_{\mu}=\pm 1$. Moreover, we remark that
@@
$\hat{C}_{\mu}(\alpha_{\mu},\beta_{\mu})|k_{\mu},l_{\mu}\rangle=k_{\mu}l_{\mu}|k_{\mu},l_{\mu}\rangle,\quad(k_{\mu},l_{\mu}=\pm 1).$ (118)
@@ A.2 Four experiments
Consider now the four independent experiments $E_{1}-E_{4}$ together. Each of the four sources emits a pair of photons in the state $|\psi_{\mu}\rangle$, so that the overall four-pair state is given by
@@
Next, using (117) we define the “Bell operator” $\hat{T}$, as
@@
[…]
@@ B.2 Calculation of $p_{k_{\mu}l_{\mu}}^{\mathrm{QM}}(\alpha_{\mu},\beta_{\mu})$
$=\frac{1}{2}\langle k_{\mu}|\biggl{[}|e_{+}^{(A_{\mu})}\rangle\!\langle e_{+}^{(A_{\mu})}|+|e_{-}^{(A_{\mu})}\rangle\!\langle e_{-}^{(A_{\mu})}|\biggr{]}|k_{\mu}\rangle$
@@
Using this equation, it is easy to obtain the marginal distribution (B.2) as,
@@
Appendix D Random variable representation of the Bell operator $\hat{T}$
@@
$\phantom{=}-\cos[2(\alpha^{\prime}-\beta)]+\cos[2(\alpha^{\prime}-\beta^{\prime})].$ (150)

### v3

@@ V.1 Actual upper bound of $S_{N}$
Let us analyze the problem from a practical point of view, considering only the measured data at our disposal [29, 30, 31, 32, 2]. To measure $S_{N}$ we must perform $N$ times the series of four experiments $\{E_{\mu}\}$, that is we must generate $4N$ distinct photon pairs. After we had ran the four experiments $N$ times, at the end of the day we got an array of $4N$ values, say $V_{N}$, given by
@@
By definition, the elements of the $N\times 4$ matrix $V_{N}^{\text{HV}}$ are equal to either $+1/N$ or $-1/N$, and given by the corresponding elements of $V_{N}$ in (88). If we sum the $\mu^{\text{th}}$ column of $V_{N}^{\text{HV}}$, we obtain the correlation function (84). Clearly, rearranging the order of the elements in each column will not alter the value of the correlation function. So, we can exploit this freedom to recast $V_{N}^{\text{HV}}$ in the new matrix $W_{N}^{\text{HV}}$, defined by
@@ VI Conclusions
In summary, in this paper we have demonstrated two different things. First, no counterfactual reasoning is required to prove the CHSH inequality. Second, the upper bound of this inequality can be greater than $2$, but the numerical value of the actual bound is not computable because of our unavoidable ignorance about the probability distribution of the hidden variables. To be more specific, the Bell parameter $S$ is bounded by $\left|S\right|\leq 2$ for HV theories, where $S$ is the mean value of a linear combination of correlation functions:
@@
One final note: given the experimental data set collected, for example, in the form of the matrix $V_{N}$ in (88), a researcher can, of course, perform a full statistical analysis of these data and also infer the rate of convergence of the limit $S_{N}\to S$. However, the moment he or she wants to compare these data with a given physical model, he or she is faced with an obvious imbalance. In fact, while QM provides a complete model that makes it possible to calculate not only the mean values of certain measurable quantities but also their probability distributions, generic hidden-variable theories provide, in practice, only mean values, which are insufficient to infer all the statistical properties of the hidden-variable model. As a matter of fact, the symmetry between quantum and hidden-variable theories of our statistical analysis presented in Sec. IV, breaks down in the moment the explicit knowledge of the correlation functions $C(\alpha,\beta)$ becomes necessary.
So, in the end, what the researcher can practically do is just to verify the agreement of the experimental data with the predictions of quantum mechanics, as after all is done in all meaningful quantum physics experiments (except for those checking Bell’s theorem!) [33]. Because of the relevance of Bell’s theorem in practical applications of quantum mechanics, such as quantum computing or quantum cryptography, we believe that these results have the potential to open up new avenues on this important topic.
@@ A.1 Single experiment
where $k_{\mu},l_{\mu}=\pm 1$. Moreover,
@@
$\hat{C}_{\mu}(\alpha_{\mu},\beta_{\mu})|k_{\mu},l_{\mu}\rangle=k_{\mu}l_{\mu}|k_{\mu},l_{\mu}\rangle,$ (118)
with $k_{\mu},l_{\mu}=\pm 1$.
@@ A.2 Four experiments
Consider now the four independent experiments $E_{1},E_{2},E_{3}$ and $E_{4}$ together. Each of the four sources emits a pair of photons in the state $|\psi_{\mu}\rangle$, so that the overall four-pair state is given by [10],
@@
Next, using (117) we define the “Bell operator” $\hat{T}\in\mathcal{H}$, as
@@
$\phantom{=.}+\cos[4(\alpha^{\prime}-\beta)]+\cos[4(\alpha^{\prime}-\beta^{\prime})]\Bigr{\}},$ (131)
@@ B.2 Calculation of $p_{k_{\mu}l_{\mu}}^{\mathrm{QM}}(\alpha_{\mu},\beta_{\mu})$
$=\frac{1}{2}\langle k_{\mu}|\underbrace{\biggl{[}|e_{+}^{(A_{\mu})}\rangle\!\langle e_{+}^{(A_{\mu})}|+|e_{-}^{(A_{\mu})}\rangle\!\langle e_{-}^{(A_{\mu})}|\biggr{]}}_{=\;\hat{I}_{\mu}}|k_{\mu}\rangle$
@@
Using this equation, it is easy to obtain again the marginal distribution (B.2) as,
@@
Appendix D Random variable representation of the quantum Bell operator $\hat{T}$
@@
$\phantom{=.}-\cos[2(\alpha^{\prime}-\beta)]+\cos[2(\alpha^{\prime}-\beta^{\prime})].$ (150)
@@
D.1 Four-dimensional Hilbert space representation of the operator $\hat{T}$
The Bell operator $\hat{T}$ in (A.2), is defined in the 256-dimensional Hilbert space $\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes\mathcal{H}_{3}\otimes\mathcal{H}_{4}$, because quantum mechanics forbids the separate and simultaneous measurement of the four correlation functions in the expression (III.1) of $S(\alpha,\beta,\alpha^{\prime},\beta^{\prime})$. However, it is not really necessary to measure such functions separately and simultaneously. This kind of situation occurs quite frequently in quantum mechanics. For example, one can measure the discrete energy spectrum of a harmonic oscillator without measuring separately and simultaneously both the kinetic and potential energy, which would be impossible to do because position and momentum operators do not commute. In principle, it would not be impossible to design an experiment whose outcomes are the eigenvalues of the whole operator $\hat{S}\in\mathcal{H}^{(A)}\otimes\mathcal{H}^{(B)}$, defined by
$\hat{S}$ :${\,\vcentcolon=\,}\hat{Z}^{(A)}(\alpha)\otimes\hat{Z}^{(B)}(\beta)+\hat{Z}^{(A)}(\alpha)\otimes\hat{Z}^{(B)}(\beta^{\prime})$
$\phantom{=}+\hat{Z}^{(A)}(\alpha^{\prime})\otimes\hat{Z}^{(B)}(\beta)-\hat{Z}^{(A)}(\alpha^{\prime})\otimes\hat{Z}^{(B)}(\beta^{\prime}).$ (153)
From (D.1) it follows that $\hat{S}$ is defined in the 4-dimensional two-photon Hilbert space. In this equation
where
$\begin{split}|\zeta_{+}^{(F)}(\phi)\rangle&=\cos\phi|e_{+}^{(F)}\rangle+\sin\phi|e_{-}^{(F)}\rangle,\\[6.0pt] |\zeta_{-}^{(F)}(\phi)\rangle&=-\sin\phi|e_{+}^{(F)}\rangle+\cos\phi|e_{-}^{(F)}\rangle,\\[6.0pt] \end{split}$ (155)
with $|\zeta_{\pm}^{(F)}(0)\rangle=|e_{\pm}^{(F)}\rangle$, and $(F,\phi)=(A,\alpha)$, or $(F,\phi)=(B_{,}\beta)$.
The operator $\hat{S}$ is self-adjoint by construction, therefore it could legitimately represent an actual observable. Using von Neumann’s spectral theorem [34, 35], it is not difficult to calculate the probability density function $f_{S}^{|\Psi\rangle}(s)$ for the random variable $S$ associated with the operator $\hat{S}$ and the state vector $|\Psi\rangle$. As usual, this density is defined by
$f_{S}^{|\Psi\rangle}(s)=\langle\Psi|\delta\bigl{(}s\,\hat{I}^{(A)}\otimes\hat{I}^{(B)}-\hat{S}\bigr{)}|\Psi\rangle.$ (156)
A straightforward calculation gives
$f_{S}^{|\Psi\rangle}(s)$ $=\frac{1}{2}\left(1+\frac{E}{s_{0}}\right)\delta(s-s_{0})$
$\phantom{=}+\frac{1}{2}\left(1-\frac{E}{s_{0}}\right)\delta(s+s_{0}),$ (157)
where we have defined
$s_{0}=2\sqrt{1-\sin\left[2(\alpha-\alpha^{\prime})\right]\sin\left[2(\beta-\beta^{\prime})\right]},$ (158)
and
$E$ $=-\cos[2(\alpha-\beta)]-\cos[2(\alpha-\beta^{\prime})]-\cos[2(\alpha^{\prime}-\beta)]$
$\phantom{.x.}+\cos[2(\alpha^{\prime}-\beta^{\prime})].$ (159)
Note that (D.1) is correct because although $\hat{S}$ has four distinct eigenvalues ±$\pm s_{0},\pm s_{1}$, the eigenvectors associated with ±$\pm s_{1}$ are orthogonal to $|\Psi\rangle$, so that they do not contribute to $f_{S}^{|\Psi\rangle}(s)$. Then, from (D.1) it follows that
$\langle\Psi|\hat{S}|\Psi\rangle$ $=\int_{\mathbb{R}}s\,f_{S}^{|\Psi\rangle}(s)\,\textrm{d}s$
$=E,$ (160)
as it should be.
So, things work fine with the operator $\hat{S}$. However, the problem is that while every physical observable of a quantum system is represented by a self-adjoint operator, the vice versa we do not know whether it is true or not [38]. Therefore, until someone is able to mount an experiment whose outcomes are the eigenvalues of the operator $\hat{S}$, we cannot claim that $\hat{S}$ represents a physical observable. We remark that there has recently been a proposal to make these four measurements at once [39] based on weak value measurements of incompatible observables [40, 41].